What have been the percentage change in the value of Bitcoin expressed in U.S. dollars over the l...

  1. What have been the percentage change in the value of Bitcoin expressed in U.S. dollars over the last 30 days? What has been the one-year change and three-year change in the value of Bitcoin? Do you detect a trend?
  2. Is Bitcoin money? Does Bitcoin fulfill the functions of money? If you conclude that Bitcoin is not money what is it? Explain.
Answer #1
  1. The value of Bitcoin expressed in USD as of 10th April 2019 is 5304.8 and that on March 10,2019 was 4000.
  • The %change in value of Bitcoin= (Value on April - Value on March)/Value on March *100

= (5304.8-4000)/4000 * 100

= 1304.8/4000 *100

= .3262*100 = 32.62%

2. The value of Bitcoin as of April 2019 is $5312.19 and that as of April 2018 was $9240.0

  • The one year %change in value of Bitcoin= (5312.19-9240)/9240* 100

= -3927.81/9240 *100

=-.4250 *100 =42.508%(decrease)

3. Value of bitcoin as of April 2016 was $438.9.

  • The three year %change in value of Bitcoin= (5312.19- 438.9)/438.9 *100   

= (4873.29)/438.9 *100

=11.103 *100 = 1,110.34%

  • The trend that is visible from the above information provided is that there has been a huge change in the value of Bitcoin since the last three years and that the daily/ monthly basis change trends also show the same. However, an unexplained factor does arise when we see that the value of Bitcoin has rather fallen in the last one year.
  • The three basic functions of money are that is serves as a medium of exchange,as a store of value and a unit of account.  Although Bitcoins are used as a medium of exchange at the world level by a lot of people, there function as store of value and unit of account are not definite due to the volatility attached with their transaction and value. Thus they cannot be wholly classified as money as they do not fulfill the basic function of money. Thus we can say that Bitcoin is not money. Even if we say that Bitcoin is money we are sure that they are no legal money. Also to add Bitcoin is money to people on internet who trade in it.
  • Bitcoin is a from of electronic cash more commonly known as cryptocurrency. It is an digital currency without legal authority that can be shared from person to person without need of intermediaries.